At what pressure will 36 grams of CO2 occupy 28.9 L at standard temperature


1
Expert's answer
2021-05-31T02:02:00-0400

MM(CO2) = 44 g/mol

n(CO2) "= \frac{36}{44}=0.818 \;mol"

V=28.9 L

T= 273 K

Ideal Gas Law:

pV=nRT

R = 0.08206 L×atm/mol×K

"p= \frac{nRT}{V} \\

= \frac{0.818 \times 0.08206 \times 273}{28.9} = 0.6342 \;atm"

Answer: 0.6342 atm